In how many years will it take a certain investment to double itself if the interest is compounded annually at 6%?

Answer by Alan3354(69443)

In how many years will it take a certain investment to double itself if the interest is compounded annually at 6%?
----------------
FV = PV*(1.06)^t
(1.06)^t = 2
t*log(1.06) = log(2)
t = log(2)/log(1.06)